OpenStudy (anonymous):

how many photons of light having a wavelength 4000 armstrong are necessary to provide 1.00J of energy?

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@Frostbite

OpenStudy (anonymous):

I think you mean angstroms. So first you have to find the frequency of the photons \[frequency=C/\] 3.00*10^8 / 4*10^-7= 7.5*10^14 Hz Then you can find the energy of one photon Energy=h*frequency So 6.62*10^-34 * 7.5*10^14= 5.0*10^-19 J So then 1.00J / 5.0*10^-19= 2.0*10^18 photons

OpenStudy (anonymous):

frequency=C/lambda

OpenStudy (anonymous):

4000 angstroms is equal to 4*10^-7 M